please help me solve this problem . . thank you
The distance between two automobiles is 140 km. If the cars are driven toward each other, they will meet in 48 minutes. However, if they are driven in the same direction, they will meet in 4 hours. What is the rate at which each car is driven?

Answer by amoresroy(361)

Let x = the speed of faster car
y = the speed of slower car
Cars are driven towards each other:
(48/60) (x+y) = 140
x + y = 175
Cars are driven in same direction:
4y + 140 = 4x
4y = 4x-140
y = x-35
Substitute y=x-35 in 1st equation to get
x + (x-35) = 175
2x = 210
x = 105
y = 105-35 = 70
The cars are driven at the rate of 105 and 70 kph.